OpenStudy (anonymous):

In a circle, a 90° sector has area 36π ft2. What is the radius of the circle? A. 18 ft B. 16 ft C. 6 ft D. 12 ft

jimthompson5910 (jim_thompson5910):

a 90° sector takes up 90/360 = 1/4 of a full circle so if the area of this sector is 36pi square feet, then the total area of the circle is 4*36pi = 144pi square feet

jimthompson5910 (jim_thompson5910):

what this means is that we can use this to find the radius A = pi*r^2 144pi = pi*r^2 now solve for r

OpenStudy (anonymous):

12?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you